OpenStudy (anonymous):

Help Pleaseee!(: Been stuck on it for an hr now. Give the acceleration of a=d^2s/dt^2, initial velocity, and initial position of a body moving on a coordinate line. Find the body's position at time t, When a=9.8, v(0)=-3, and s(0)=0.

OpenStudy (anonymous):

How long have you been suck on it for?

OpenStudy (anonymous):

An hour now.. and I have no idea how to go about doing this problem lol

OpenStudy (anonymous):

the joke was you said suck instead of stuck , and then I asked you again using the word suck in place of stuck

OpenStudy (anonymous):

ohhhh... lol okay fixing that now..

OpenStudy (anonymous):

s=ut+1/2at^2

OpenStudy (anonymous):

I don't get that that function represents? @Krishnadas

OpenStudy (anonymous):

ok ...i will tell you @pdd21

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

ok..i dont know if this helps..but this is one way to the answwer

OpenStudy (anonymous):

have you studies that equation in Physics?

OpenStudy (anonymous):

*studied

OpenStudy (anonymous):

@pdd21

OpenStudy (anonymous):

you here?

OpenStudy (anonymous):

@pdd21 need you here dude

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

reply......:(

OpenStudy (anonymous):

sorry... got distracted trying the problem out again.. and i haven't studied the equation in physics.

OpenStudy (anonymous):

ok i will help

OpenStudy (anonymous):

let acceleration=a,u=initial velocity,and V =velocity at time t..okay?

OpenStudy (anonymous):

a=dv/dt

OpenStudy (anonymous):

dv=adt

OpenStudy (anonymous):

integrate both sides \[\int\limits_{u}^{v}dv=\int\limits_{0}^{t}a dt\]

OpenStudy (anonymous):

I havent really learned how to integrate yet, so it would be weird if I used it to show my work on this problem lol

OpenStudy (anonymous):

that will be a problem..But i can show you how we got that equation

OpenStudy (anonymous):

okay..(:

OpenStudy (anonymous):

aint there a way to anti derive? its the same as integrating right?

OpenStudy (anonymous):

ya..integration=finding anti derivate

OpenStudy (anonymous):

so when time changes from 0 to t,velocity changes from u to V..you got this idea?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

then you know nothing abt integration??

OpenStudy (anonymous):

yup! lol... I'm pretty much having a hard time with this stuff lol

OpenStudy (anonymous):